# Conflicts in Texts: Data, Implications and Challenges

### v1

@@ Abstract
As NLP models become increasingly integrated into real-world applications, it becomes clear that there is a need to address the fact that models often rely on and generate conflicting information. Conflicts could reflect the complexity of situations, changes that need to be explained and dealt with, difficulties in data annotation, and mistakes in generated outputs. In all cases, disregarding the conflicts in data could result in undesired behaviors of models and undermine NLP models’ reliability and trustworthiness. This survey categorizes these conflicts into three key areas: (1) natural texts on the web, where factual inconsistencies, subjective biases, and multiple perspectives introduce contradictions; (2) human-annotated data, where annotator disagreements, mistakes, and societal biases impact model training; and (3) model interactions, where hallucinations and knowledge conflicts emerge during deployment. While prior work has addressed some of these conflicts in isolation, we unify them under the broader concept of conflicting information, analyze their implications, and discuss mitigation strategies. We highlight key challenges and future directions for developing conflict-aware NLP systems that can reason over and reconcile conflicting information more effectively.
@@ 1 Introduction
Figure 1: Examples of the three different ares of conflicts discussed in this work. The first example describes a case where two different entities of the same name are found naturally on the web, the second example elaborates the annotation disagreement in a sentiment analysis task, and the third showcases a knowledge conflict between the context and memory of LLMs during model interactions.
@@
for tree= grow=east, parent anchor=east, child anchor=west, edge=draw, thick, rounded corners, draw=black, fit=band, minimum width=5cm, align=center, anchor=west, s sep=18pt, l sep=22pt, inner sep=10pt [Conflicts in NLP, fill=yellow!40, tier=0 [Model Interactions, fill=yellow!30, tier=1 [Hallucination, fill=green!40, tier=2 [Contextual Hallucination, fill=red!40, tier=3 [Maynez et al. (2020),
Raunak et al. (2021), Dale et al. (2023), fill=cyan!40] ] [Factual Hallucination, fill=red!40, tier=3 [Lin et al. (2022),
Pagnoni et al. (2021), Honovich et al. (2021), fill=cyan!40] ] ] [Knowledge Conflicts, fill=green!40, tier=2 [Across Different Models, fill=red!40, tier=3 [Cohen et al. (2023),
Zhu et al. (2024), Zhao et al. (2024), fill=cyan!40] ] [Between Context and Parameters, fill=red!40, tier=3 [Longpre et al. (2021); Chen et al. (2022a),
Chen et al. (2021a); Lazaridou et al. (2021), fill=cyan!40] ] ] ] [Human Annotations, fill=yellow!30, tier=1 [Ethical and Societal Biases, fill=green!40, tier=2 [Annotation Biases, fill=red!40, tier=3 [Sap et al. (2022), Faisal et al. (2022),
Thorn Jakobsen et al. (2022) , fill=cyan!40] ] ] [Human Judgements, fill=green!40, tier=2 [Interpretation Disagreements, fill=red!40, tier=3 [Kahneman (2021), Uma et al. (2021),
Sandri et al. (2023), Jiang and de Marneffe (2022), fill=cyan!40] ] ] ] [Natural Conflicts on Web, fill=yellow!30, tier=1 [Conflicts in Opinions, fill=green!40, tier=2 [Framing Bias, fill=red!40, tier=3 [Card et al. (2015), Liu et al. (2019),
Fan et al. (2019), Lei et al. (2022), fill=cyan!40] ] [Perspectives, fill=red!40, tier=3 [Chen et al. (2019), Liu et al. (2021a),
Lee et al. (2022) , Wan et al. (2024) , fill=cyan!40] ] ] [Factual Conflicts, fill=green!40, tier=2 [Contradictory Evidence, fill=red!40, tier=3 [Chen et al. (2022a),Hong et al. (2024),
Liu et al. (2024), Pham et al. (2024), fill=cyan!40] ] [Ambiguity, fill=red!40, tier=3 [Min et al. (2020), Zhang and Choi (2021),
Dhingra et al. (2022), Cole et al. (2023), fill=cyan!40] ] ] ] ]
@@
Existing work on conflicts in NLP tends to focus on specific issues, such as annotation disagreements Uma et al. (2021), hallucinations and factuality Zhang et al. (2023); Wang et al. (2023), and knowledge conflicts Xu et al. (2024); Feng et al. (2024), without synthesizing these problems into a broader perspective. In this survey, we conceptualize these diverse challenges under the umbrella of conflicting information and analyze their origins, implications, and mitigation strategies. We first examine conflicts inherent in training data, including natural texts from the web and human-annotated datasets. We then explore conflicts arising during interactions with models in the LLM era, discussing their impact on downstream tasks and highlighting key challenges and future directions for conflict-aware AI systems.
The abundance of online data is accompanied by inherent conflicts, stemming from diverse sources, interpretations, and biases. These conflicts manifest as factual conflicts, such as semantic ambiguities Pavlick and Tetreault (2016); Min et al. (2020) and factual inconsistencies Pham et al. (2024); Liu et al. (2024), or as conflicts in opinions related to political ideologies Entman (1993); Recasens and et al. (2013) and perspectives Chen et al. (2019); Liu et al. (2021a). Factual conflicts are particularly prevalent in open-domain question answering (QA) and retrieval-augmented generation (RAG) systems Chen et al. (2017), where aggregating knowledge from multiple sources introduces inconsistencies Liu et al. (2024). These challenges highlight the need for conflict-aware retrieval and reasoning mechanisms to improve model reliability Xie et al. (2024). Unlike factual conflicts, opinionated disagreements reflect the variability in human interpretation, beliefs, and ideological stances Chen et al. (2019); Fan et al. (2019). The presence of conflicting viewpoints complicates tasks such as summarization, sentiment analysis, and dialogue generation, where maintaining coherence and neutrality is crucial Liu et al. (2021a); Lee et al. (2022). Furthermore, the uneven distribution and biases of web data also affects models to behave from a Western perspective Ramaswamy et al. (2023); Mihalcea et al. (2024).
@@
By systematically categorizing and analyzing these conflicts, this survey provides a unified perspective on their origins, implications, and mitigation strategies. Addressing these issues is essential for building robust and trustworthy AI systems that operate effectively across diverse domains and user groups. Our findings contribute to the development of conflict-aware frameworks for data collection, model training, and model usage, ultimately enhancing the fairness and reliability of NLP.
@@ 2.1.1 Origins
Min et al. (2020) was the first work to study the effects of ambiguity in open domain question answering. They introduced AmbigQA, a dataset highlighting that over half of the open-domain, natural questions are ambiguous, with diverse sources of ambiguity such as event and entity references. Zhang and Choi (2021) proposed the SituatedQA task, showing that a significant fraction of open-domain questions are valid only under particular temporal or geographic contexts. Many other work specifically focus on the temporal aspect of ambiguity, benchmarking and evaluating models’ awareness and adaptation to time-sensitive questions Chen et al. (2021b); Liska et al. (2022); Kasai et al. (2023).
@@
Conflicts in NLP systems arise when information on the web presents conflicting evidence towards a factual question. This issue is particularly prevalent in open-domain question answering settings, where models must navigate inconsistencies across diverse information sources. Liu et al. (2024) find that 25% of unambiguous factual questions queried on Google retrieve conflicting evidence from multiple sources. For instance, a Google search for "When was Kendrick Lamar’s first album released?" yields conflicting evidence, illustrating the challenge of integrating contradictory information in retrieval-based QA systems Liu et al. (2024).
Researchers have proposed different datasets to systematically study how NLP models handle such conflicts. Li et al. (2024b) introduce ContraDoc, a human-annotated dataset of long documents with internal contradictions; Pham et al. (2024) propose WhoQA, a benchmark dataset that constructs conflicts by formulating questions about a shared property among entities with the same name (e.g. "Who is George Washington?"); and Liu et al. (2024) construct QACC, a human-annotated dataset of conflicting results retrieved by Google. Beyond empirical datasets, several studies have proposed synthetic approaches to simulate conflicts through entity substitution Chen et al. (2022a); Hong et al. (2024), machine-generated conflicting evidence Pan et al. (2023); Wan et al. (2024); Hong et al. (2024), and pre-defined rule-based templates Kazemi et al. (2023).
@@ 2.1.2 Implications and Mitigation
Zhang and Choi (2021) show that pre-trained language models perform competitively at identifying whether a question is context dependent, matching human agreements, but lag behind human-level performance by a significant margin when answering questions dependent on temporal contexts. Cole et al. (2023) demonstrate that large language models benefit from a “disambiguate-then-answer” pipeline, showing improved reliability by detecting ambiguity before attempting an answer. Dhingra et al. (2022) propose time-aware language models that condition on timestamps to mitigate confusion arising from outdated facts or evolving knowledge.
Studies have shown that retrieval performance heavily impacts which sources models rely on Chen et al. (2022a), knowledge conflicts within contexts significantly degrade LLMs’ performance in RAG settings Pham et al. (2024); Liu et al. (2024); Li et al. (2024b), and QA models are vulnerable to even small amounts of evidence contamination brought by misinformation Pan et al. (2023). In addition, large language models (LLMs) exhibit a strong confirmation bias, favoring external evidence that aligns with their parametric memory, even when conflicting evidence is present Xie et al. (2024). To mitigate these issues, Hong et al. (2024) proposed fine-tuning a discriminator or prompting GPT-3.5 to elicit its discriminative capability and show that these approaches significantly enhance model robustness. On the other hand, Liu et al. (2024) proposed fine-tuning LLMs with human-written explanations to teach models to reason through conflicting evidence.
@@ 2.2.1 Origins
Individuals and communities often hold diverse perspectives on the same issue. Such diversity is evident in online discussions and debates, where the multiplicity of viewpoints can lead to conflicting opinions. For instance, on controversial topics such as "Animals should have lawful rights," people express varying stances Chen et al. (2019), posing challenges for downstream tasks like summarization where consolidating viewpoints and presenting unbiased information are crucial Liu et al. (2021a); Lee et al. (2022).
Chen et al. (2019) address this challenge by presenting a range of perspectives on a given claim. The authors introduce the task of substantiated perspective discovery, where a system identifies diverse, evidence-supported perspectives that take a stance on a claim, and curated a dataset, PERSPECTRUM, for the task, using online debate platforms and search engines. Wan et al. (2024) introduce ConflictingQA, a dataset comprising controversial questions paired with real-world evidence documents presenting divergent facts, argumentation styles, and conclusions. Plepi et al. (2024) study perspective-taking in contentious online conversations (e.g. social media dilemmas). The authors create a new corpus of 95k conflict scenarios augmented with each user’s self-disclosed background information. Liu et al. (2021a) propose MultiOpEd, an open-domain corpus focusing on automatic perspective discovery. MultiOpEd comprises 1,397 controversial topics, each accompanied by two editorials expressing opposing viewpoints, and each editorial includes a one-sentence perspective summarizing its core argument and a brief abstract highlighting supporting details.
@@
Various studies have examined different facets of such media bias. Card et al. (2015) introduce the Media Frames Corpus (MFC) to facilitate the computational study of media framing in news articles. The MFC comprises several thousand news articles covering three policy issues, each annotated according to 15 general-purpose framing dimensions. Liu et al. (2019) introduce Gun Violence Frame Corpus (GVFC), a dataset of news headlines with frames curated and annotated by journalism and communication experts. Fan et al. (2019) focus on informational bias, where factual content is subtly framed by content selection and organization within articles. They introduce BASIL, a dataset comprising 300 news articles annotated with 1,727 bias spans, revealing that informational bias occurs more frequently than lexical bias. Lei et al. (2022) identify ideological bias at the sentence level by analyzing news discourse structure, showing that bias-indicative sentences may appear neutral in isolation.
@@ 2.2.2 Implications and Mitigation
Analysis of PERSPECTRUM reveals significant natural language understanding challenges, as human performance substantially outperforms machine baselines at identifying diverse, evidence-supported perspectives Chen et al. (2019). Furthermore, when selecting real-world evidence for controversial questions, LLMs predominantly prioritize the relevance of the evidence to the query, often disregarding stylistic attributes such as the presence of scientific references or a neutral tone Wan et al. (2024). In addition, the distribution and biases of web data also affects models to behave from a Western perspective Ramaswamy et al. (2023); Mihalcea et al. (2024). Studies have shown that LLMs’ outputs skew toward the values of Western English-speaking countries Tao et al. (2024); Naous et al. (2024), and misalignment is more pronounced for underrepresented personas and on culturally sensitive topics such as social values Al Kuwatly et al. (2020). Furthermore, LLMs often provide inconsistent answers to the same question when prompted in different languages Li et al. (2024a); AlKhamissi et al. (2024), revealing conflicting cultural perspectives within a single model.
Several studies have proposed methods to address conflicts in perspectives. Liu et al. (2021a) show that incorporating auxiliary tasks enhances the quality of perspective summarization. Chen et al. (2022b) propose a novel document retrieval paradigm that focuses aggregating and displaying responses from web documents based on varying viewpoints, and reveal that users prefer seeing search results in different clusters of perspectives instead of in a list ranked by relevance. Jiang et al. (2023) tackle opinion summarization for user reviews by generating summaries from different perspectives. Their framework selects subsets of reviews based on sentiment polarity and informational contrast, and produces balanced pros, cons, and verdict summaries. Plepi et al. (2024) demonstrate that a tailored generation model which conditions on a user’s personal context produces more appropriate and empathetic responses than large general models, effectively capturing different viewpoints in the conflict.
To mitigate framing bias and ideology conflicts, Milbauer et al. (2021) propose a method to uncover complex ideological and worldview differences across online communities beyond a single left-vs-right axis. Their study identifies multiple axes of polarization and nuanced ideological distinctions, offering a more multifaceted analysis of online opinion differences. Liu et al. (2022) introduce a pre-training approach for ideology classification that learns ideological bias by directly comparing news articles about the same event reported by outlets with different leanings. Chen et al. (2023) tackle political ideology classification under limited and biased data by disentangling content from style, enabling accurate ideology detection with minimal training data. Lee et al. (2022) present a model that leverages news titles and employs hierarchical multi-task learning to neutralize biased content from title to article, while Liu et al. (2023) induce a neutral event graph that captures events with minimal framing bias by synthesizing across different ideologies.

### v2

@@ Abstract
As NLP models become increasingly integrated into real-world applications, it becomes clear that there is a need to address the fact that models often rely on and generate conflicting information. Conflicts could reflect the complexity of situations, changes that need to be explained and dealt with, difficulties in data annotation, and mistakes in generated outputs. In all cases, disregarding the conflicts in data could result in undesired behaviors of models and undermine NLP models’ reliability and trustworthiness. This survey categorizes these conflicts into three key areas: (1) natural texts on the web, where factual inconsistencies, subjective biases, and multiple perspectives introduce contradictions; (2) human-annotated data, where annotator disagreements, mistakes, and societal biases impact model training; and (3) model interactions, where hallucinations and knowledge conflicts emerge during deployment. While prior work has addressed some of these conflicts in isolation, we unify them under the broader concept of conflicting information, analyze their implications, and discuss mitigation strategies. We highlight key challenges for developing conflict-aware and robust NLP systems, and propose concrete research directions to address them.
@@ 1 Introduction
Figure 1: Examples of the three different areas of conflicts discussed in this work. The first example describes a case where two different entities of the same name are found naturally on the web, the second example elaborates the annotation disagreement in a sentiment analysis task, and the third showcases a knowledge conflict between the context and memory of LLMs during model interactions.
Existing work on conflicts in NLP tends to focus on specific issues, such as annotation disagreements Uma et al. (2021); Klie et al. (2023), hallucinations and factuality Zhang et al. (2023); Wang et al. (2023), and knowledge conflicts Xu et al. (2024); Feng et al. (2024), without synthesizing these problems into a broader perspective. In this survey, we conceptualize these diverse challenges under the umbrella of conflicting information and analyze their origins, implications, and mitigation strategies.
@@
t
@@
To ensure comprehensive and representative coverage of conflicts in NLP, we first establish a high-level categorization encompassing three primary sources: (1) natural conflicts present in web data, (2) conflicts arising from human annotation, and (3) conflicts emerging from model interactions. Notably, conflicts found in natural web texts and human-annotated datasets are primarily present in the training data—i.e., the inputs to models—whereas conflicts involving model interactions can arise in various forms, such as inconsistencies between model outputs and their inputs, contradictions among multiple outputs, or conflicts within the outputs themselves. For each category, we identify influential and widely cited survey papers as initial seed works  Uma et al. (2021); Klie et al. (2023); Zhang et al. (2023); Xu et al. (2024); Feng et al. (2024); Wang et al. (2023). Building upon these seeds, we systematically trace and incorporate the most impactful and representative studies for each type of conflict through citation chaining and targeted literature searches across major databases. This approach enables us to synthesize developments in each category and connect them, thereby providing an integrated discussion of current challenges, impacts on downstream tasks, and promising future directions for conflict-aware AI systems (See Appendix B for more discussion about our survey methodology).
The abundance of online data is accompanied by inherent conflicts, stemming from diverse sources, interpretations, and biases. These conflicts manifest as factual conflicts, such as semantic ambiguities Pavlick and Tetreault (2016); Min et al. (2020) [^1] The previous version on ACL Anthology has a wrong paper title for a reference here. and factual inconsistencies Pham et al. (2024); Liu et al. (2024), or as conflicts in opinions related to political ideologies Entman (1993); Recasens and et al. (2013) and perspectives Chen et al. (2019); Liu et al. (2021). Factual conflicts are particularly prevalent in open-domain question answering (QA) and retrieval-augmented generation (RAG) systems Chen et al. (2017), where aggregating knowledge from multiple sources introduces inconsistencies Liu et al. (2024). These challenges highlight the need for conflict-aware retrieval and reasoning mechanisms to improve model reliability Xie et al. (2024). Unlike factual conflicts, opinionated disagreements reflect the variability in human interpretation, beliefs, and ideological stances Chen et al. (2019); Fan et al. (2019). The presence of conflicting viewpoints complicates tasks such as summarization, sentiment analysis, and dialogue generation, where maintaining coherence and neutrality is crucial Liu et al. (2021); Lee et al. (2022). Furthermore, the uneven distribution and biases of web data also affects models to behave from a Western perspective Ramaswamy et al. (2023); Mihalcea et al. (2025).
@@
In this survey, we systematically examine the landscape of conflicts in NLP by categorizing them into three primary sources. For each conflict type, we detail how such conflicts arise and in what forms they take (origins), the challenges they pose (implications), and the strategies developed to address them (mitigation). We present a comprehensive taxonomy in Figure 2, as well as structured summary tables—Table 1, Table 2, and Table 3—that synthesize datasets, methodologies, and analysis from prior work. By offering a unified framework for understanding and addressing conflicting information in NLP, this survey contributes to the development of conflict-aware frameworks for data collection, model training, and model usage, ultimately enhancing the fairness and reliability of NLP.
@@ 2.1.1 Origins
Min et al. (2020) was the first work to study the effects of ambiguity in open domain question answering. They introduced AmbigQA, a dataset highlighting that over half of the open-domain, natural questions are ambiguous, with diverse sources of ambiguity such as event and entity references. Zhang and Choi (2021) proposed the SituatedQA task, showing that a significant fraction of open-domain questions are valid only under particular temporal or geographic contexts. Many other work specifically focus on the temporal aspect of ambiguity, benchmarking and evaluating models’ awareness and adaptation to time-sensitive questions Chen et al. (2021); Liska et al. (2022); Kasai et al. (2023).
@@
Conflicts in NLP systems arise when information on the web presents conflicting evidence towards a factual question. This issue is particularly prevalent in open-domain question answering settings, where models must navigate inconsistencies across diverse information sources. For example, Liu et al. (2024) find that 25% of unambiguous factual questions queried on Google retrieve conflicting evidence from multiple sources.
Researchers have proposed different datasets to systematically study how NLP models handle such conflicts. Li et al. (2024b) introduce ContraDoc, a human-annotated dataset of long documents with internal contradictions; Pham et al. (2024) propose WhoQA, a benchmark dataset that constructs conflicts by formulating questions about a shared property among entities with the same name (e.g. "Who is George Washington?"); and Liu et al. (2024) construct QACC, a human-annotated dataset of conflicting results retrieved by Google. Beyond empirical datasets, several studies have proposed synthetic approaches to simulate conflicts through entity substitution Chen et al. (2022a); Hong et al. (2024), machine-generated conflicting evidence Pan et al. (2023); Wan et al. (2024a); Hong et al. (2024), and pre-defined rule-based templates Kazemi et al. (2023).
@@ 2.1.2 Implications and Mitigation
Implications
Factual conflicts pose significant challenges for NLP systems. Pre-trained language models accurately detect context-dependent questions but fall short when answering queries requiring temporal context, performing notably below human levels Zhang and Choi (2021). Additionally, large language models (LLMs) often exhibit confirmation bias, favoring retrieved information that aligns with their parametric memory despite contradictory evidence Xie et al. (2024). Consequently, conflicting information sources severely impact retrieval-augmented generation (RAG) frameworks, significantly degrading model performance even with minimal misinformation exposure Pham et al. (2024); Liu et al. (2024); Li et al. (2024b); Pan et al. (2023).
Mitigation
To address these challenges, various mitigation strategies have been proposed. Effective methods include fine-tuning calibrators for selective abstention Chen et al. (2022a), employing a "disambiguate-then-answer" pipeline to detect ambiguity proactively Cole et al. (2023), and developing time-aware models that condition responses on timestamps to manage outdated information Dhingra et al. (2022). Further robustness improvements have been achieved through fine-tuning discriminators or prompting GPT-3.5 models to explicitly recognize conflicting evidence Hong et al. (2024), as well as incorporating human-written explanations in fine-tuning processes to enhance models’ reasoning capabilities Liu et al. (2024).
@@ 2.2.1 Origins
Individuals and communities often hold diverse perspectives on the same issue. Such diversity is evident in online discussions and debates, where the multiplicity of viewpoints can lead to conflicting opinions. For instance, on controversial topics such as "Animals should have lawful rights," people express varying stances Chen et al. (2019), posing challenges for downstream tasks like summarization where consolidating viewpoints and presenting unbiased information are crucial Liu et al. (2021); Lee et al. (2022).
Several studies have explored perspectives in the context of conflicting information. Chen et al. (2019) introduce the task of substantiated perspective discovery, where systems identify diverse, evidence-supported stances on a claim, and release the PERSPECTRUM dataset using online debates and search results. Wan et al. (2024a) propose ConflictingQA, a dataset of controversial questions paired with real-world documents that present divergent facts, arguments, and conclusions. Plepi et al. (2024) examine perspective-taking in contentious online discourse, curating a corpus of 95k conflict scenarios annotated with users’ self-reported backgrounds. Liu et al. (2021) present MultiOpEd, a corpus of 1,397 controversial topics, each paired with opposing editorials and concise summaries capturing their core perspectives.
@@
Numerous studies have investigated different aspects of media bias. Card et al. (2015) introduce the Media Frames Corpus (MFC), a collection of news articles annotated with 15 general-purpose framing dimensions across three policy issues, enabling computational analysis of media framing. Liu et al. (2019) present the Gun Violence Frame Corpus (GVFC), a dataset of news headlines annotated by domain experts to capture framing in gun violence reporting. Fan et al. (2019) examine informational bias—bias conveyed through content selection and structure—and release BASIL, a dataset of 300 news articles annotated with 1,727 bias spans, demonstrating that informational bias is more prevalent than lexical bias.
@@ 2.2.2 Implications and Mitigation
Implications
Analysis of PERSPECTRUM reveals significant natural language understanding challenges, as human performance substantially outperforms machine baselines at identifying diverse, evidence-supported perspectives Chen et al. (2019). Furthermore, when selecting real-world evidence for controversial questions, LLMs predominantly prioritize the relevance of the evidence to the query, often disregarding stylistic attributes such as the presence of scientific references or a neutral tone Wan et al. (2024a). In addition, the distribution and biases of web data also affects models to behave from a Western perspective Ramaswamy et al. (2023); Mihalcea et al. (2025). Studies have shown that LLMs’ outputs skew toward the values of Western English-speaking countries Tao et al. (2024); Naous et al. (2024), and misalignment is more pronounced for underrepresented personas and on culturally sensitive topics such as social values Al Kuwatly et al. (2020). Furthermore, LLMs often provide inconsistent answers to the same question when prompted in different languages Li et al. (2024a); AlKhamissi et al. (2024); Eloundou et al. (2025), revealing conflicting cultural perspectives within a single model.
Mitigation
Several studies have proposed methods to address conflicts in perspectives and ideological bias. Liu et al. (2021) show that auxiliary tasks improve perspective summarization quality, while Chen et al. (2022b) propose a retrieval paradigm that clusters documents by viewpoint, revealing users’ preference for diverse perspectives over relevance-ranked lists. Jiang et al. (2023b) generate opinion summaries by selecting review subsets based on sentiment polarity and contrast, producing balanced pros, cons, and verdicts. Plepi et al. (2024) demonstrate that conditioning generation on users’ personal contexts yields more empathetic and appropriate responses than general-purpose models.
To mitigate framing and ideological bias, Milbauer et al. (2021) uncover nuanced worldview differences across communities by identifying multiple axes of polarization beyond the traditional left–right spectrum. Liu et al. (2022b) pre-train models for ideology detection by comparing reporting on the same events across partisan sources. Chen et al. (2023) disentangle content from style to enable ideology classification under data scarcity and bias. Lee et al. (2022) employ hierarchical multi-task learning to neutralize bias from news titles to article bodies, while Liu et al. (2023) construct neutral event graphs by synthesizing perspectives across ideological divides.